OpenStudy (anonymous):

What is the fraction 12/28 expressed in lowest terms?

OpenStudy (susanna):

simplify it

OpenStudy (anonymous):

12/2 and 28/2 right

OpenStudy (anonymous):

i got 6/14

OpenStudy (susanna):

\[\frac{ 12 }{ 28 } = \frac{ 6 }{ 14 } = \frac{ 3 }{ 7 }\]

OpenStudy (anonymous):

ok so had to divide it again cause it wasent to the lowest term

OpenStudy (susanna):

yes

OpenStudy (anonymous):

yeahhhh